Sodium bicarbonate is commercially known as baking soda. How many grams and atoms of Oxygen are there in 15g of sodium bicarbonate?
1
Expert's answer
2020-10-06T09:16:51-0400

Solution:

Sodium bicarbonate is a chemical compound with the formula NaHCO3.


The molar mass of NaHCO3 is 84 g/mol.

The atomic mass of Oxygen is 16 g/mol.


Now, In 84 g NaHCO3 contains = (16 × 3) = 48 g Oxygen

hence, 15 g NaHCO3 contains = (48 × 15) / 84 = 8.571 g Oxygen.

Mass of Oxygen is 8.571 g.


Moles of Oxygen = Mass / Atomic mass = (8.571 g) / (16 g/mol) = 0.5357 mol

One mole of any substance contains 6.02×1023 units of that substance.

Thus:

Atoms of Oxygen = 0.5357 mol × 6.02×1023 atoms/mol = 3.225×1023 atoms of Oxygen


Answer: 8.571 grams and 3.225×1023 atoms of Oxygen are there in 15 g of sodium bicarbonate.